OpenStudy (anonymous):

Owners of a recreation area are filling a small pond with water. They are adding water at a rate of 30 liters per minute. There are 600 liters in the pond to start. Let W represent the amount of water in the pond (in liters), and let T represent the number of minutes that water has been added. Write an equation relating W to T , and then graph your equation using the axes below.

OpenStudy (anonymous):

@wio a little help?

OpenStudy (anonymous):

@dan815 could you help?

OpenStudy (anonymous):

I want you to try on your own first.

OpenStudy (anonymous):

I don't know how to do it! D: I need more of an explanation than anything else @wio

OpenStudy (anonymous):

What do you think the slope is?

OpenStudy (anonymous):

30/600 @wio

OpenStudy (anonymous):

That isn't it.

OpenStudy (anonymous):

After a minute, how many liters has it increased?

OpenStudy (anonymous):

30 so 1/30? or 30/1? @wio

OpenStudy (anonymous):

30/1

OpenStudy (anonymous):

@wio okay

OpenStudy (anonymous):

Do you remember the equation of a line, the slope intercept form?

OpenStudy (anonymous):

@wio y=mx+b?

OpenStudy (anonymous):

What is the intercept?

OpenStudy (anonymous):

600? @wio

OpenStudy (anonymous):

Yes, so what is the equation of the line?

OpenStudy (anonymous):

y=30/1+600 @wio

OpenStudy (anonymous):

What happened to \(x\)?

OpenStudy (anonymous):

By the way 30/1 = 30, they're the same thing.

OpenStudy (anonymous):

So, putting in the \(x\), what is the equation?

OpenStudy (anonymous):

y=30/1x+600 sorry I always forget the x @wio

OpenStudy (anonymous):

Okay, but don't write the "/1" you don't need it.

OpenStudy (anonymous):

whoops sorry @wio

OpenStudy (anonymous):

y=30x+600

OpenStudy (anonymous):

They want us to use W and T, so we have: \[ W=30T+600 \]

OpenStudy (anonymous):

@wio could you help me graph it?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Start out drawing your coordinate plane.

OpenStudy (anonymous):

@Mertsj can I get a mod up on here to remove this guy?^

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know how to plot a point?

OpenStudy (anonymous):

yes we can use (0.600) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is W when T is 1?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Nope, \[ W=30(1)+600 \]

OpenStudy (anonymous):

okay w=630

OpenStudy (anonymous):

So what point does that make?

OpenStudy (anonymous):

@Callisto can you help with withyourdestiny?

OpenStudy (anonymous):

(630,0)?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Nope, it is (1, 630)

OpenStudy (anonymous):

So graph that point.

OpenStudy (anonymous):

okay, @wio

OpenStudy (anonymous):

did it @wio

OpenStudy (anonymous):

Take a ruler between both points.

OpenStudy (anonymous):

Then draw a line!

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@wio